# Shrinking vs. expanding: the evolution of spatial support in degenerate Keller–Segel systems

### v1

@@ 4.2 Case of support expansion
    for all $t\in[0,T\raisebox{4.0pt}{\scriptsize$\star$}]$, and as $\partial_{ss}\overline{w}_{\mathrm{mid}}$ is continuous on the compact set $[r_{0}^{n},\rho(t)-\delta]$ for all $t\in[0,T\raisebox{4.0pt}{\scriptsize$\star$}]$, our desired regularity properties for $\overline{w}$ are immediately evident. Moreover, since $\partial_{ss}\overline{w}_{\mathrm{mid}}(s,t)<0$, we see that $\partial_{s}\overline{w}_{\mathrm{mid}}(s,t)>0$ for all $s\in[r_{0}^{n},\rho(t)-\delta)$ and $t\in[0,T\raisebox{4.0pt}{\scriptsize$\star$}]$.
@@
    by combining this with (4.31). Since $\overline{w}_{\mathrm{out}}$ is constant and hence fulfills $\mathcal{P}_{\varepsilon}\overline{w}_{\mathrm{out}}=0$, we obtain (4.29). ∎

### v2

@@ 4.2 Case of support expansion
for all $t\in[0,T\raisebox{4.0pt}{\scriptsize$\star$}]$, and as $\partial_{ss}\overline{w}_{\mathrm{mid}}$ is continuous on the compact set $r_{0}^{n},\rho(t)-\delta]$ for all $t\in[0,T\raisebox{4.0pt}{\scriptsize$\star$}]$, our desired regularity properties for $\overline{w}$ are immediately evident. Moreover, since $\partial_{ss}\overline{w}_{\mathrm{mid}}(s,t)<0$, we see that $\partial_{s}\overline{w}_{\mathrm{mid}}(s,t)>0$ for all $s\in[r_{0}^{n},\rho(t)-\delta)$ and $t\in[0,T\raisebox{4.0pt}{\scriptsize$\star$}]$.
@@
by combining this with (4.31). Since $\overline{w}_{\mathrm{out}}$ is constant and hence fulfils $\mathcal{P}_{\varepsilon}\overline{w}_{\mathrm{out}}=0$, we obtain (4.29). ∎